OpenStudy (anonymous):

If 4/9=2/x, what do you know about their cross products? Their cross products are reciprocals of each other. Their cross products are opposite of each other. Their cross products are not related. Their cross products are equal.

OpenStudy (anonymous):

The cross products are reciprocals of each other

OpenStudy (dmndlife24):

\[\frac{ 4 }{ 9 } = \frac{ 2 }{ x }\] \[9(2) = 4(x)\]

OpenStudy (dmndlife24):

^ cross products are equal

OpenStudy (anonymous):

oh true thanks